Stock A has a dividend yield of 8% but no capital gain. Stock B offers a capital gain but no dividend. If a corporate investor in the 21% tax bracket earns the same after-tax return from the two stocks, what capital gain does B offer?

Multiple Choice

  • 8.29%

  • 9.06%

  • 6.00%

  • 11.31%
Answer #1

Answer is option B

B. 9.06%

RA= 0.08 − (0.08 × 0.50 × 0.21) = 0.0716, or 7.16%

RB = 0.0716 / (1 − 0.21) = 0.906, or 9.06%